OpenStudy (anonymous):

Imagine you are teaching a fellow student how to solve: 4 + 2(x − 1) = 5x − (x + 10) In your own words, explain the process for solving this equation. Please include your work at each step of the process along with the final answer.

OpenStudy (anonymous):

im from colorado...i will brb im going out for lunch today, if you have any more questions just mention them to me, thnx cody

OpenStudy (anonymous):

4 + 2(x − 1) = 5x − (x + 10) 4 + 2x - 2 = 5x - x - 10 2x + 2 = 4x - 10 2=2x - 10 12 = 2x 6 = x do the rest on your own